1. (a) [15 marks] Total ascorbic acid can be analysed by spectrophotometric method based on the oxidation of ascorbic acid toI dont know how to do question 1 part (iv)
Answer #1

Dear student kindly find the complete solution of Part (1V) OF question 1(a) and give your positive feedback if you properly understood the solution..

if any problem let me know...

Warm Regards

Dr. Rakhee

PART (iv) LOD ☺ LOB (Limit of Blank) = mean blank + 1.645 (SD blank) = 0.0071 +1-645 (0.0009) ☺ = 0.0071 +0.0014805 = 0.00858